6.0cm length measured with a ruler marked in mm what is the absolute uncer


1
Expert's answer
2020-10-09T07:03:11-0400

The absolute uncertainty of a measurement device is equal to the half of its the smalles mark. If the ruler is marked in mm, then the absolute uncertainty of this measurement is:


"\Delta l =\dfrac{1mm}{2}= 0.5mm = 0.5\times 10^{-3}m"

Answer. "\Delta l =0.5\times 10^{-3}m".